OpenStudy (anonymous):

What is the angle of rotation for a regular nonagon? (Note: Be sure to type out the word degrees, instead of the symbol.)

OpenStudy (anonymous):

360/no of sides = 360/9 = 40 degree

OpenStudy (anonymous):

once again thank you !

OpenStudy (anonymous):

once again welcome :)